OpenStudy (anonymous):

Multiply and simplify: (4t − 9)2 16t2 + 81 16t2 − 72t + 81 16t2 − 72t − 81 16t2 − 81

OpenStudy (anonymous):

Use the identity \((a-b)^2 = a^2 -2ab + b^2\) Put a= 4t and b=9 into it. \((4t-9)^2 = (4t)^2 -2(4t)(9) + 9^2 = ...?\)

OpenStudy (jiteshmeghwal9):

(4t-9)^2 =>(4t)^2+(-9)^2+2(4a)(-9)=??

OpenStudy (jiteshmeghwal9):

=16t^2+81-72t

OpenStudy (jiteshmeghwal9):

clear????

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

It's clear because you've given the asker the answer.

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (jiteshmeghwal9):

yw:)

OpenStudy (jiteshmeghwal9):

LOL @RolyPoly